# Chirality-induced selectivity of angular momentum by orbital Edelstein effect in carbon nanotubes

### v1

@@ Results
Lastly, we would like to note that this strict chirality dependence of the effect only is applicable in this energy range near the Fermi level. Fig. 5b shows the same plot at an energy of $E=7.91\,\mathrm{eV}$ where, depending on the tube, we see vastly different band structures and multiple states with different $L_{\nu,z}(k)$ exist. As a result, the blue data points are scattered and even the $(5,2)$ and $(10,4)$ nanotubes that are characterized by the same chirality angle have strongly different $\chi_{z,\mathrm{atom}}^{L_{z}}$, even with different signs. The values are $1.55\times 10^{-12}\,\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$ versus $-4.48\times 10^{-12}\,\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$, respectively.
Discussion
@@
Near the Fermi level, the effect is purely governed by the chirality of the structure and scales linearly with the circumference of the nanotube. However, note that this means that the orbital Edelstein susceptibility per cross-sectional area $A$ scales inversely with the circumference, $m_{L_{z}}/A\propto\frac{1}{C}\sin(\alpha/6)\mathcal{E}_{z}$. For this reason, growing many small nanotubes in a finite area, can result in a larger CIOS than having only a few large nanotubes, as long as the nanotubes can be positioned closely together. It might also be interesting to consider multi-walled CNTs where multiple nanotubes can be positioned inside of each other, thereby increasing or decreasing the effect, depending on the respective chiralities.
@@
Finding an approximate relation between the orbital angular momentum $L_{z}$ in CNTs and the velocity $v_{C}$ in graphene along $\boldsymbol{C}$ (as presented before) is also an important result because it allows us to connect our findings with previous calculations from the 1990s. In fact, before the modern formulation of orbital magnetization was established [63, 64] and before CISS has been observed [16], a ‘chiral conductivity’ had been theoretically discussed that is calculated via $v_{C}$ of the two-dimensional material forming the tube; Ref. [67] for BC2N nanotubes and Ref. [68] for BN nanotubes. This chiral conductivity has even been related to self-inductance [69] meaning that a chiral current gives rise to magnetic properties. Therefore, the chiral conductivity is closely related to the orbital Edelstein susceptibility that we calculate here giving rise to CIOS. Indeed, Ref. [70] finds similar oscillating energy-dependent curves of the chiral conductivity but the authors only considered insulating CNTs.
@@
Our results show that the orbital angular momentum couples to the chirality of a system and exists even without spin-orbit coupling. In semi-classical terms, the origin of CIOS is in real-space as there are helical hopping paths along chiral CNTs. However, in comparison to CIOS in a simple helix [61], as in tellurium or selenium, the Dirac physics and sub-band structure in CNTs have an important influence on the effect. They make a quantum mechanical treatment crucial, while in Ref. [61], the effect was perfectly analogous to a classical motion along a helix. In comparison to a conventional Rashba system, where an Edelstein effect occurs only at an interface with the magnetization perpendicular to the current [59, 60], here we generate a homogeneous magnetization all over the CNT along the current direction.
@@
In an experiment, the orbital magnetic moment generated by CIOS can be detected as a spin moment leading to CISS, due to a conversion in the electrodes caused by their spin-orbit coupling. In future studies, related chiral systems like nanotubes formed by multiple elements, chiral nano-carbons, chiral carbon-based heterostructures or molecular-functionalized nanotubes can be analyzed, for some of which CISS has been observed experimentally already [24, 25, 26, 27, 28, 29, 30]. The efficient and controllable filtering of spin and orbital angular momentum allows for the generation of spin and orbital torques. This makes chiral CNTs promising candidates for applications in quantum computing and spin-orbitronics.
@@ Methods
The tight-binding Hamiltonian consists of a nearest-neighbor hopping term for the radial $p_{z}$ orbitals of the carbon atoms
@@
Here, $c_{i}^{\dagger}$ and $c_{j}$ are the creation and annihilation operators of an electron at site $i$ and $j$, respectively. The bracket indicates nearest neighbors. In our example of a $(n,m)=(7,1)$ nanotube, the unit cell consists of $76$ atoms, each with 3 nearest neighbors. $t=2.8\,\mathrm{eV}$ is the hopping amplitude.
@@
We consider periodic copies of the tube along $x$ and $y$ but do not consider any coupling between them. This allows us to use $k_{x}$ and $k_{y}$ in the Hamiltonian in order to calculate the $z$ component of the orbital angular momentum according to the modern formulation of orbital magnetization [62, 63, 64, 65, 66]
$L_{\nu,z}(k)=i\frac{m_{e}}{g_{L}\hbar}\sum_{\mu\neq\nu}\frac{1}{E_{\mu}(k)-E_{\nu}(k)}$ (3)
×$\times\left[\langle\nu k|\nicefrac{{\partial H}}{{\partial k_{x}}}|\mu k\rangle\langle\mu k|\nicefrac{{\partial H}}{{\partial k_{y}}}|\nu k\rangle-(\nu\leftrightarrow\mu)\right].$
Here, $\nu$ is the band index with $\lvert nk\rangle$ the eigenvector and $E_{\nu}(k)$ the eigenenergy determined by diagonalizing the Hamiltonian $H$. $m_{e}$ is the electron mass, $g_{L}=1$ is the Landé g-factor and $\hbar$ the reduced Planck constant. Note that the trick of using the super cell is equivalent to defining $L_{\nu,z}$ by means of a real-space tight-binding Hamiltonian. This formulation includes inter-site contributions that are missing in the atomic-center approximation, which is often used in the literature.
@@
The data that support the findings of this work are available at DOI: [Added upon publication].
@@ Acknowledgements
This work was supported by the EIC Pathfinder OPEN grant 101129641 “Orbital Engineering for Innovative Electronics”. B.G. performed calculations, prepared the figures, planned the project and wrote the manuscript with significant inputs from all authors. All authors discussed the results.
@@
accompanies this paper at [insert link].

### v2

@@ Abstract
Lastly, we would like to note that this strict chirality dependence of the effect only is applicable in this energy range near the Fermi level. Fig. 5b shows the same plot at an energy of $E=7.91\,\mathrm{eV}$ where, depending on the tube, we see vastly different band structures and multiple states with different $L_{\nu,z}(k)$ exist. As a result, the blue data points are scattered and even the $(5,2)$ and $(10,4)$ nanotubes that are characterized by the same chirality angle have strongly different $\chi_{z,\mathrm{atom}}^{L_{z}}$, even with different signs. The values are $1.55\times 10^{-12}\,\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$ versus $4.48\times 10^{-12}\,\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$, respectively.
Relation to experiments and CISS
In an experiment, the orbital magnetic moment generated by CIOS could be detected as a spin moment leading to CISS, due to a conversion in the electrodes caused by their spin-orbit coupling [53]. This means that the measured signal could be affected and even dominated by CIOS which is caused directly by the chirality of the sample and is not dependent on the often small spin-orbit coupling in a chiral structure.
However, note that a feature in transport experiments predicted by this approach requires further investigation [53]: When using magnetic leads, the II-$V$ dependence is predicted to become asymmetric due to the current-induced orbital magnetic moment. However, this asymmetry is not observed in most transport experiments, although an asymmetric signature is visible in Refs. [72, 73]. The orbital mechanism discussed in this paper is therefore likely not the only origin of CISS in chiral materials but is a complementary effect to the spin-driven effect that originates from spin-orbit coupling of the chiral material itself. From the theory side, it would be interesting to quantify corrections to the orbital Edelstein effect and to calculate the spin Edelstein effect caused by curvature-induced spin-orbit coupling [74, 75] and the hybridization of different orbitals. For large radii, we expect these corrections to be negligible. Experimentally, it would be helpful to vary the spin-orbit coupling of the lead (e. g. by changing the material) and to analyze the effect on the CISS signal in transport experiments, as previously suggested [76].
When a magnetic field $B$ is applied along the CNT, it couples to the orbital angular momentum and affects the band structure accordingly [77, 78, 79]. In a unidirectional magnetoresistance measurement, the resistance $R$ depends on the sign of $B$ which gives rise to the phenomenon of electrical magnetochiral anisotropy (EMChA) that is a signature of chiral materials [80, 81, 82]. We expect that this effect occurs for chiral CNTs as well but the precise field dependence and the interplay with the electric field that generates the non-equilibrium orbital magnetic moment has to be analyzed carefully in future studies.
Besides the aforementioned potential for spin-to-charge interconversion, chiral CNTs could also be used directly for orbital-to-charge interconversion. The experimental observation of orbital magnetic moments and orbital currents has progressed rapidly over the past few years, as summarized in the review paper [83]. Detection has been achieved: (i) By magneto-optical Kerr effect measurements [41, 42], (ii) by detecting charge currents corresponding to inverse orbital effects [55], (iii) by measuring the Hanle magnetoresistance [84], (iv) via measuring orbital torques [85]. Especially noteworthy for our findings is a previous measurement of the inverse orbital Edelstein effect in a two-dimensional electron gas at an oxide interface [55]. Based on our calculations of the orbital Edelstein susceptibility, chiral CNTs act as filters and polarizers for orbital currents which could potentially be detected by the mentioned techniques and mechanisms.
Furthermore, CIOS could be measured more easily by photoemission than in transport experiments [53]: Chiral molecules play the same role as spiral phase plates for photoelectrons and their total detectable magnetic moment includes the dominating orbital angular momentum. This means that the spin moment, that may come from the spin-orbit coupling of the sample or substrate, would not interfere with the orbital contribution as much as in transport experiments.
Conclusion
@@
Near the Fermi level, the effect is purely governed by the chirality of the structure and scales linearly with the circumference of the nanotube. However, note that this means that the orbital Edelstein susceptibility per cross-sectional area $A$ scales inversely with the circumference, $m_{L_{z}}/A\propto\frac{1}{C}\sin(6\alpha)\mathcal{E}_{z}$. For this reason, growing many small nanotubes in a finite area, can result in a larger CIOS than having only a few large nanotubes, as long as the nanotubes can be positioned closely together. It might also be interesting to consider multi-walled CNTs where multiple nanotubes can be positioned inside of each other, thereby increasing or decreasing the effect, depending on the respective chiralities.
@@
Finding an approximate relation between the orbital angular momentum $L_{z}$ in CNTs and the velocity $v_{C}$ in graphene along $\mathbf{C}$ (as presented before) is also an important result because it allows us to connect our findings with previous calculations from the 1990s. In fact, before the modern formulation of orbital magnetization was established [64, 65] and before CISS has been observed [16], a ‘chiral conductivity’ had been theoretically discussed that is calculated via $v_{C}$ of the two-dimensional material forming the tube (for BC2N nanotubes [86] and for BN nanotubes [87]). This chiral conductivity has even been related to self-inductance [88] meaning that a chiral current gives rise to magnetic properties. Therefore, the chiral conductivity is closely related to the orbital Edelstein susceptibility that we calculate here giving rise to CIOS. Indeed, similar oscillating energy-dependent curves of the chiral conductivity have been calculated [89] but the authors only considered insulating CNTs.
@@
Our results show that the orbital angular momentum couples to the chirality of a system and exists even without spin-orbit coupling. In semi-classical terms, the origin of CIOS is in real-space as there are helical hopping paths along chiral CNTs. However, in comparison to CIOS in a simple helix [62], as in tellurium or selenium, the Dirac physics and sub-band structure in CNTs have an important influence on the effect. They make a quantum mechanical treatment crucial, while in a toy model of a helix [62] the effect was perfectly analogous to a classical motion along a helix. In comparison to a conventional Rashba system, where an Edelstein effect occurs only at an interface with the magnetization perpendicular to the current [59, 60], here we generate a homogeneous magnetization all over the CNT along the current direction.
@@
In future studies, related chiral systems like nanotubes formed by multiple elements, chiral nano-carbons, chiral carbon-based heterostructures or molecular-functionalized nanotubes can be analyzed, for some of which CISS has been observed experimentally already [24, 25, 26, 27, 28, 29, 30]. The efficient and controllable filtering of spin and orbital angular momentum allows for the generation of spin and orbital torques. This makes chiral CNTs promising candidates for applications in quantum computing and spin-orbitronics.
@@ Methods
The tight-binding Hamiltonian considers the radial $p_{z}$ orbitals of the carbon atoms
@@
Other orbitals are not required to resemble ab initio calculations, especially for large tube radii [90]. Here, $c_{i}^{\dagger}$ and $c_{j}$ are the creation and annihilation operators of an electron at site $i$ and $j$, respectively. The bracket indicates nearest neighbors. In our example of a $(n,m)=(7,1)$ nanotube, the unit cell consists of $76$ atoms, each with 3 nearest neighbors. $t=2.8\,\mathrm{eV}$ is the hopping amplitude.
We consider periodic copies of the tube along $x$ and $y$ but do not consider any coupling between them. Therefore, even though the band structure exhibits no dispersion along $k_{x}$ and $k_{y}$, the Hamiltonian depends on these inplane components of the wave vector which accounts for the chirality of the nanotube. The individual hopping terms from atom $j$ to $i$ have a phase $\mathbf{k}\cdot\mathbf{r}_{ij}$. Here $\mathbf{r}_{ij}$ is the connecting vector that has in general $x$, $y$ and $z$ components. By using this convention, the eigenvectors correspond to the lattice-periodic part of the Bloch wave from which the orbital angular momentum can be calculated directly.
@@
We calculate the $z$ component of the orbital angular momentum according to the modern formulation of orbital magnetization [63, 64, 65, 66, 67]
$L_{\nu,z}(k)=\mathrm{i}\frac{m_{e}}{g_{L}\hbar}\sum_{\mu\neq\nu}\frac{1}{E_{\mu}(k)-E_{\nu}(k)}$ (3)
$\times\left[\langle\nu k|\nicefrac{{\partial H}}{{\partial k_{x}}}|\mu k\rangle\langle\mu k|\nicefrac{{\partial H}}{{\partial k_{y}}}|\nu k\rangle-(\nu\leftrightarrow\mu)\right].$
Here, $\nu$ is the band index with $\lvert\nu k\rangle$ the eigenvector and $E_{\nu}(k)$ the eigenenergy determined by diagonalizing the Hamiltonian $H$. Note that the eigenvectors correspond to the lattice-periodic part of the Bloch wave. $m_{e}$ is the electron mass, $g_{L}=1$ is the Landé g-factor and $\hbar$ the reduced Planck constant. Note that the trick of using the super cell is equivalent to defining $L_{\nu,z}$ by means of a real-space tight-binding Hamiltonian. This formulation includes inter-site contributions that are missing in the atomic-center approximation, which is often used in the literature.
@@
The data that support the findings of this work are available at DOI: 10.5281/zenodo.17065253.
@@ Acknowledgements
B.G. performed calculations, prepared the figures, planned the project and wrote the manuscript with significant inputs from I.M. and S.L.. B.G., I.M. and S.L. discussed the results.
Additional information
@@
The online version contains supplementary material available at [insert link].